Distinguish btw stationary and progressive waves


1
Expert's answer
2021-02-23T14:12:25-0500
  1. Stationary waves do not transfer energy from one location to another while progressive waves transfers energy from one place to another excluding transfer of matter.
  2. The equation of a stationary wave is: "y=Asin\frac{2\pi }{\lambda }\left(vt\right)" while the equation of a progressive wave is: "y=Asin\frac{2\pi }{\lambda }\left(vt-x\right)".
  3. In stationary waves, only other particles except ones at the nodes performs periodic motion while in progressive waves, all particles of the medium implements periodic motion.
  4. In stationary waves, amplitudes of particles between successive nodes are different while in progressive waves, amplitudes of any two particles which are in phase are the same